OpenStudy (luigi0210):

Use log differentiation:

OpenStudy (luigi0210):

\[\Huge y=(ln x)^{cos7x}\]

OpenStudy (luigi0210):

So far.. \[\LARGE lny=cosx~ln(lnx)\] Now differentiation, and product rule..?

OpenStudy (luigi0210):

@ganeshie8 Right so far? And that's suppose to be cos7x

ganeshie8 (ganeshie8):

two ln's may not be needed

OpenStudy (whpalmer4):

You'll end up with enough logs to build a cabin :-)

OpenStudy (luigi0210):

Here's the full question.. http://prntscr.com/30chur

OpenStudy (agent0smith):

\[\Large \ln x^a = a \ln x\]\[\Large (\ln x)^a \ne a \ln x\]

ganeshie8 (ganeshie8):

ahh yes i see, cos^7x is to the whole log... so yes luigi is rigiht

OpenStudy (agent0smith):

\[\LARGE lny=cosx~\ln(lnx) \]just use the product and chain rule now

OpenStudy (luigi0210):

So I end up with.. \[\LARGE \frac{y'}{y}=cos7x(\frac{1}{lnx}*\frac{1}{x})+-7sin7x(ln(lnx))\]?

ganeshie8 (ganeshie8):

looks good, and the answer in ur snap looks right... oly thing is the parenthesis are missing everywhere

ganeshie8 (ganeshie8):

once u fix them, i think it wud show checkmark :)

OpenStudy (luigi0210):

Omg, you were right, it was the parenthesis ._. Thanks everyone, for everything :)